# Bayesian Monocular Depth Refinement via Neural Radiance Fields

### v1

@@ Abstract
Monocular depth estimation has applications in many fields, such as autonomous navigation and extended reality, making it an essential computer vision task. However, current methods often produce smooth depth maps that lack the fine geometric detail needed for accurate scene understanding. We propose MDENeRF, an iterative framework that refines monocular depth estimates using depth information from Neural Radiance Fields (NeRFs). MDENeRF consists of three components: (1) an initial monocular estimate for global structure, (2) a NeRF trained on perturbed viewpoints, with per-pixel uncertainty, and (3) Bayesian fusion of the noisy monocular and NeRF depths. We derive NeRF uncertainty from the volume rendering process to iteratively inject high-frequency fine details. Meanwhile, our monocular prior maintains global structure. We demonstrate superior performance on key metrics and experiments using indoor scenes from the SUN RGB-D dataset.
@@ I Introduction
We propose MDENeRF, an iterative depth refinement framework that improves our initial monocular depth estimate. First, we synthetically create nearby viewpoints using perturbed images. These simulate a multi-view environment that enhances NeRF training. The depth learned from the NeRF can be reprojected to the original pose, and we statistically formulate the per-pixel uncertainty of the resulting NeRF depth map. Both our monocular estimate and our trained NeRFs are noisy, so we use Bayesian inference to fuse the two estimates based on their uncertainty. We find that this preserves global structure during refinement while iteratively refining high-frequency details without drift. Extensive quantitative and qualitative evaluations demonstrate robust depth refinement. Code is available at https://github.com/ArunMut/MDENeRF.
@@ III-F Iterative Refinement Loop
Each iteration outputs a refined depth $D_{\text{refined}}$ and uncertainty $\sigma_{\text{refined}}^{2}$, which initialize the next cycle. Drift is avoided because Bayesian fusion anchors refinement to the monocular prior and injects NeRF detail only where confidence is high, reverting elsewhere. Thus, our iterative refinement has minimal error accumulation; in practice, 2–3 iterations suffice.
@@ IV Experiments
Monocular Depth Baseline and Setup. We use MiDaS (DPT-Large) for the initial monocular estimate $D_{0}$ in all our experiments [17]. Since MiDaS produces scale-ambiguous depth, ground-truth depth is accessed only during evaluation via median scale alignment. The actual refinement steps of MDENeRF are scale-free and do not access ground-truth. We perform evaluations on 100 randomly selected indoor scenes on the SUN RGB-D dataset [18].
Implementation and Metrics. We use small pose perturbations, generating N=10N{=}10 synthetic views per scene. This preserves the original $640{\times}480$ resolution with minimal artifacts ($2\%$ hole-filling). Within each scene, the NeRF is trained for 2000 optimization steps and refined for two iterations, as additional iterations result in diminishing returns. After median scale alignment, we report the global accuracy with mean squared error (MSE). High-frequency details are evaluated using edge sharpness (mean depth gradient magnitude, normalized relative to MiDaS) and edge F1. Additionally, we analyze our uncertainty quality with the correlation between predicted uncertainty and true depth error. For numerical stability, we clamp extremely small variances and skip affine calibration in degenerate cases, reverting to the monocular estimate.
@@ IV-A Quantitative Results
Table I shows that MDENeRF significantly improves high-frequency depth structures while maintaining strong global accuracy. Compared to the MiDaS baseline, MDENeRF increases edge sharpness by $18\%$ and improves edge F1 by $6.3$ points. Together, these metrics indicate sharper and more accurate depth boundaries. Furthermore, global depth error is reduced overall, with MDENeRF achieving the lowest MSE among all methods, demonstrating that the overall scene layout does not degrade. Ultimately, these results highlight the effectiveness of our fusion algorithm in refining MDE.
TABLE I: Average performance on SUN RGB-D test scenes. Global errors are reported after median scale alignment. Edge metrics are normalized relative to MiDaS.
@@
Figure 2 confirms that MDENeRF’s uncertainty estimates are well calibrated and meaningful. After grouping pixels from low to high predicted uncertainty, it plots the mean absolute depth error (MAE) as a function of predicted uncertainty percentile. Depth error increases monotonically with uncertainty, meaning regions MDENeRF labels with higher uncertainty are consistently less accurate. This trend is supported by strong rank correlations (Spearman $\rho=0.968$, Kendall $\tau=0.905$) and a high Pearson correlation ($r=0.860$), indicating that predicted uncertainty reliably reflects true depth error. The standard error of our estimates increases with higher uncertainty, but this is expected behavior due to smaller sample sizes and inherently ambiguous regions that lead to higher error variability.
@@ IV-C Ablation Studies
Our fusion mechanism is closed-form and does not contain hyperparameters during inference. Therefore, we ablate our modeling choices rather than tuning parameters.
@@
We make three main observations from our ablation study. (i) Uncertainty: replacing NeRF variance with a constant reduces edge sharpness and degrades uncertainty-aware behavior, indicating that ray termination variance is a meaningful confidence signal. (ii) Precision-weighted multi-view aggregation: replacing our proposed fusion with heuristic minimum aggregation modestly degrades both global and local quality, especially in regions that are affected by disocclusion. (iii) Calibration: without the affine WLS alignment, global accuracy drops noticeably, indicating that the scale and shift must be corrected prior to probabilistic fusion. Finally, edge sharpness did slightly improve when removing the monocular prior, but global error significantly worsened, confirming its role in stabilizing refinement and preventing drift.
@@ V Conclusions
We present MDENeRF, a framework that iteratively refines monocular depth estimates via Bayesian fusion with NeRF depth. We interpret NeRF volume rendering weights as ray termination distributions, allowing for per-pixel depth uncertainty estimation. Fusing NeRF and monocular depth using Bayesian inference allows MDENeRF to selectively enhance high-frequency details without degrading the global scene structure. Our indoor experiments validate MDENeRF’s fidelity in improving local accuracy, preserving global information, and calibrating meaningful uncertainty.
Current limitations of MDENeRF include scalability to larger scenes and more complex geometry. Furthermore, our NeRF training incurs computational costs that may be mitigated in future work. Potential directions involve multi-scale NeRFs and frequency-based analysis for more targeted refinement. Dynamic scene support is an open challenge, and such parallel lines of research may greatly support monocular depth refinement. These advances in depth estimation will ultimately benefit safety-critical computer vision applications.

### v2

@@ Abstract
Monocular depth estimation has applications in many fields, such as autonomous navigation and extended reality, making it an essential computer vision task. However, current methods often produce smooth depth maps that lack the fine geometric detail needed for accurate scene understanding. We propose MDENeRF, an iterative framework that refines monocular depth estimates using depth information from Neural Radiance Fields (NeRFs). MDENeRF consists of three components: (1) an initial monocular estimate for global structure, (2) a NeRF trained on perturbed viewpoints, with per-pixel uncertainty, and (3) Bayesian fusion of the noisy monocular and NeRF depths. We derive NeRF uncertainty from the volume rendering process to iteratively inject high-frequency fine details. Meanwhile, our monocular prior maintains global structure. We demonstrate improvements on key metrics and experiments using indoor scenes from the SUN RGB-D dataset.
footnotetext: © 2026 IEEE. Accepted for publication in the 2025 8th International Conference on Algorithms, Computing and Artificial Intelligence (ACAI). Personal use is permitted. All other uses require permission from IEEE.
@@ I Introduction
We propose MDENeRF, an iterative depth refinement framework that improves our initial monocular depth estimate (See Figure 1). First, we synthetically create nearby viewpoints using perturbed images. These simulate a multi-view environment that enhances NeRF training. The depth learned from the NeRF can be reprojected to the original pose, and we statistically formulate the per-pixel uncertainty of the resulting NeRF depth map. Both our monocular estimate and our trained NeRFs are noisy, so we use Bayesian inference to fuse the two estimates based on their uncertainty. We find that this preserves global structure during refinement while iteratively refining high-frequency details. Quantitative and qualitative evaluations demonstrate robust depth refinement. Code is available at https://github.com/ArunMut/MDENeRF.
@@ III-F Iterative Refinement Loop
Each iteration outputs a refined depth $D_{\text{refined}}$ and uncertainty $\sigma_{\text{refined}}^{2}$, which initialize the next cycle. Bayesian fusion anchors refinement to the monocular prior and injects NeRF detail only where confidence is high, reverting elsewhere. Thus, our iterative refinement has minimal error accumulation; in practice, 2–3 iterations suffice.
@@ IV Experiments
Monocular Depth Baseline and Setup. We use MiDaS (DPT-Large) for the initial monocular estimate $D_{0}$ in all our experiments [16]. Since MiDaS produces scale-ambiguous depth, ground-truth depth is accessed only during evaluation via median scale alignment. The actual refinement steps of MDENeRF do not use ground-truth scale or metric supervision. We perform evaluations on 20 randomly selected indoor scenes on the SUN RGB-D dataset [17].
Implementation and Metrics. We use small pose perturbations, generating N=10N{=}10 synthetic views per scene. This preserves the original $640{\times}480$ resolution with minimal artifacts ($2\%$ hole-filling). Within each scene, the NeRF is trained for 2000 optimization steps and refined for two iterations, as additional iterations result in diminishing returns. Although we use a low-fidelity NeRF, MDENeRF is plug-and-play and can improve from stronger NeRFs in future work. After median scale alignment, we report the global accuracy with mean squared error (MSE). High-frequency details are evaluated using edge sharpness (mean depth gradient magnitude, normalized relative to MiDaS) and edge F1. Additionally, we analyze our uncertainty quality with the correlation between predicted uncertainty and true depth error. For numerical stability, we clamp extremely small variances and skip affine calibration in degenerate cases, reverting to the monocular estimate.
@@ IV-A Quantitative Results
Table I shows that MDENeRF consistently improves high-frequency depth structures while maintaining strong global accuracy. Compared to the MiDaS baseline, MDENeRF increases edge sharpness by $9\%$ and improves edge F1 by a relative amount of $2.9\%$. Together, these metrics indicate sharper and more accurate depth boundaries. Furthermore, global depth error is relatively maintained overall, with MDENeRF degrading by $1.92\%$ in MSE, indicating global error is largely preserved despite sharper boundaries. Ultimately, these results highlight the effectiveness of our fusion algorithm in refining MDE.
TABLE I: Average performance on SUN RGB-D test scenes. Global errors are reported after median scale alignment. Edge sharpness is normalized relative to MiDaS.
@@
Figure 2 supports the formulations behind MDENeRF’s uncertainty estimates. After grouping pixels from low to high predicted uncertainty, it plots the mean absolute depth error (MAE) as a function of predicted uncertainty percentile. In the low-to-mid uncertainty regime, depth error increases with predicted uncertainty, indicating that regions MDENeRF labels as more uncertain tend to be less accurate. However, this trend does not hold across the entire uncertainty spectrum (Spearman $\rho=-0.038$), suggesting that predicted uncertainty is less informative in high-uncertainty regions. We attribute this behavior to the limited capacity of our lightweight NeRF model, which can trivially be improved in future work with stronger NeRFs.
@@ IV-C Ablation Studies
Our fusion mechanism is closed-form and does not contain fusion parameters during inference. Therefore, we ablate our modeling choices rather than tuning parameters.
@@
As shown in Table II, we make three main observations after ablating. (i) Uncertainty: replacing NeRF variance with a constant reduces edge sharpness and degrades uncertainty-aware behavior, indicating that ray termination variance is a meaningful confidence signal. (ii) Precision-weighted multi-view aggregation: replacing our proposed fusion with heuristic minimum aggregation modestly degrades both global and local quality, especially in regions that are affected by disocclusion. (iii) Calibration: without the affine WLS alignment, global accuracy drops noticeably, indicating that the scale and shift must be corrected prior to probabilistic fusion. Finally, edge sharpness did slightly improve when removing the monocular prior, but global error significantly worsened, confirming its role in stabilizing refinement.
@@ V Conclusions
We present MDENeRF, a framework that iteratively refines monocular depth estimates via Bayesian fusion with NeRF depth. We interpret NeRF volume rendering weights as ray termination distributions, allowing for per-pixel depth uncertainty estimation. Fusing NeRF and monocular depth using Bayesian inference allows MDENeRF to selectively enhance high-frequency details while preserving the global scene structure. Our indoor experiments validate MDENeRF’s fidelity in improving local accuracy, preserving global information, and calibrating uncertainty.
Current limitations of MDENeRF include scalability to larger scenes and more complex geometry. Furthermore, our NeRF training incurs computational costs that may be mitigated in future work. Potential directions involve multi-scale NeRFs and frequency-based analysis for more targeted refinement. Dynamic scene support is an open challenge, and such parallel lines of research may greatly support monocular depth refinement. Future work can readily integrate our plug-and-play framework into higher-fidelity NeRFs for improved uncertainty calibration and depth refinement. These advances in depth estimation will ultimately benefit safety-critical computer vision applications.